# Not Eliminate but Aggregate: Post-Hoc Control over Mixture-of-Experts to Address Shortcut Shifts in Natural Language Understanding

### v1

@@ 1 Introduction
A simple solution to this problem is to eliminate reliance on shortcuts, which is the mainstream approach, including recent studies in NLU (Clark et al., 2019; He et al., 2019; Mahabadi et al., 2020). Typically, those methods up-weight training instances where some known shortcuts cannot predict labels correctly and down-weight the others. A practical deficiency of this approach arises in a performance trade-off between ID and OOD data. It deviates models from ID data by eliminating shortcuts, which are valid features in ID data. Due to this trade-off, another practical problem arises where the hyperparameter search has to be made using OOD test or validation data, as noticed as the limitation of previous work (Clark et al., 2019; Mahabadi et al., 2020; Clark et al., 2020a; Ghaddar et al., 2021; Liu et al., 2021; Creager et al., 2021; Yu et al., 2022; Yang et al., 2023). Even when using OOD development data, its distribution is the same as that of test data. Thus, in other words, the approach requires knowing the test-time distribution to tune hyperparameters, which is impractical in testing OOD robustness.
@@ 2.3 Problems of Previous Approaches
Initial studies used training data where shortcuts are pre-identified, in addition to OOD test or validation data (Clark et al., 2019; Arjovsky et al., 2019; Sagawa et al., 2020, inter alia). Pre-identification of shortcuts is costly as it requires careful analysis of given data. Seeking more practical solutions, subsequent approaches followed that did not require pre-identification of shortcuts in training data (Clark et al., 2020a; Liu et al., 2021; Creager et al., 2021, inter alia). However, they still need OOD test or validation data to tune hyperparameters. This requirement has been discussed as a serious limitation for practical use (Clark et al., 2019; Mahabadi et al., 2020; Clark et al., 2020a; Ghaddar et al., 2021; Liu et al., 2021; Creager et al., 2021; Yu et al., 2022). Moreover, the performance of those approaches on OOD data is considerably low without the hyperparameter tuning on OOD data (Yang et al., 2023).
@@ 3.1 Training Phase: Mixture Model to Capture Latent Features
We use this penalty term with the following modifications. First, the penalty cannot be minimized to zero when the mini-batch size exceeds the number of experts . During joint minimization with the classification loss , forcing the minimization of the never-zero penalty could lead models too far away from the optimal solution for . To avoid this, we consider that in each row of , its corresponding expert captures the same latent features among the top- elements (instances). We then exclude those elements to allow such multi-instance assignments. Let be a function to drop out the elements with the top- values in each row to 0 (Algorithm 1). We minimize .
@@ 4 Experiments
Our goal is to achieve predictions robust to the distribution shifts regarding shortcuts. In this section, we test whether the proposed post-hoc control improves performance on those OOD tests and analyze the mechanism based on our assumption.
@@ 4.3 Analyses
We analyzed the contribution to the performance with respect to the hyperparameters of our mixture model: the number of experts , the number of row-wise dropouts , and the loss-weighting value . Table 4 shows the ablation study on MNLI. This table shows how performance changes by varying one of the hyperparameters from the values determined to be optimal on the ID dev. There is little to no difference in performance on the ID dev for any value, but the OOD performance with post-hoc control is best for nearly all the values determined to be best on the ID dev. It is also worth noting that using our and top- dropout consistently improves OOD performance better than without using them (when or is zero). These results indicate the effectiveness of the proposed training strategy and hyperparameter search.
@@ 5 Related Work
Several approaches used counterfactual framework of causal inference. To make counterfactual predictions unaffected by shortcuts, Tian et al. (2022); Niu et al. (2021) combined predictions of a main model and a shortcut-dependent model. Wang and Culotta (2020) classified features into genuine or spurious and selected genuine features for predictions. Others utilized identified spurious features to train model predictions to be invariant to interventions on the spurious features (Veitch et al., 2021; Makar et al., 2022; Puli et al., 2022b).
@@
However, these methods still require OOD test data regarding shortcuts to tune hyperparameters, as described in Section 2.3. Our method is different from them in that the training and tuning can be conducted solely on ID data. In Section 4.2, we demonstrated that in the setting of fully unknown shortcuts where only ID data are available, our method improves the performance on OOD data significantly better than the previous methods.
@@
Outside of NLU, shortcuts have been addressed in the ML literature as one of the broader OOD problems (Krueger et al., 2021; Yang et al., 2023). Still, many methods used in ML and NLU tasks have the same concepts in common, such as reweighting (Nam et al., 2020; Liu et al., 2021; Clark et al., 2020a; Utama et al., 2020b), IRM (Creager et al., 2021; Yu et al., 2022), counterfactual invariance (Veitch et al., 2021; Makar et al., 2022; Puli et al., 2022b), and data augmentation (Yao et al., 2022; Puli et al., 2022a; Wu et al., 2022). As described in Section 2.2, IRM (Arjovsky et al., 2019) and GroupDRO (Sagawa et al., 2020) are the two principal approaches. These approaches considered the known shortcut setting, and similar to NLU literature, their follow-up approaches have sought to address shortcuts in the unknown shortcut setting (Nam et al., 2020; Liu et al., 2021; Creager et al., 2021; Yao et al., 2022; Puli et al., 2022a; Izmailov et al., 2022; Kirichenko et al., 2023). However, also similar to NLU literature, those follow-up approaches still require shortcuts to be pre-identified in validation sets (Yang et al., 2023).
@@ Appendix A Additional Figures
Appendix B Further Details of Setup
Table 5 specifies the URLs of the datasets, pre-trained models, and codes of the previous methods we introduced in Section 4.1.
@@ Appendix B Further Details of Setup
Datasets MNLI https://cims.nyu.edu/~sbowman/multinli/ HANS https://github.com/tommccoy1/hans QQP and PAWS https://github.com/google-research-datasets/paws FEVER and Symm. v1/v2 https://github.com/TalSchuster/FeverSymmetric Pre-Trained Models BERT https://huggingface.co/bert-base-uncased DeBERTa https://huggingface.co/microsoft/deberta-v3-large Codes Conf-reg ∗ https://github.com/UKPLab/emnlp2020-debiasing-unknown JTT∗ https://github.com/YyzHarry/SubpopBench RISK https://github.com/CuteyThyme/RISK EIIL https://github.com/PluviophileYU/BAI BAI https://github.com/PluviophileYU/BAI GroupDRO∗ https://github.com/YyzHarry/SubpopBench ReWeightCRT∗ https://github.com/YyzHarry/SubpopBench

### v3

@@ 1 Introduction
A simple solution to this problem is to eliminate reliance on shortcuts, which is the mainstream approach, including recent studies in NLU (Clark et al., 2019; He et al., 2019; Mahabadi et al., 2020). Typically, those methods up-weight training instances where some known shortcuts cannot predict labels correctly and down-weight the others. A practical deficiency of this approach arises in a performance trade-off between ID and OOD data. It deviates models from ID data by eliminating shortcuts, which are valid features in ID data. Due to this trade-off, another practical problem arises where the hyperparameter search has to be made using OOD test or validation data, as noticed as the limitation of previous work (Clark et al., 2019; Mahabadi et al., 2020; Clark et al., 2020a; Ghaddar et al., 2021; Liu et al., 2021; Creager et al., 2021; Yu et al., 2022; Yang et al., 2023). Even when using OOD validation data, its distribution is the same as that of test data. Thus, in other words, the approach requires knowing the test-time distribution to tune hyperparameters, which is impractical in testing OOD robustness.
@@ 2.3 Problems of Previous Approaches
Initial studies used training data where shortcuts are pre-identified, in addition to OOD test or validation data (Clark et al., 2019; Arjovsky et al., 2019; Sagawa et al., 2020, inter alia). Pre-identification of shortcuts is costly as it requires careful analysis of given data. Seeking more practical solutions, subsequent approaches followed that did not require pre-identification of shortcuts in training data (Clark et al., 2020a; Liu et al., 2021; Creager et al., 2021, inter alia). However, they still need OOD test or validation data related to the pre-identified shortcuts to tune hyperparameters. This requirement has been discussed as a serious limitation for practical use (Clark et al., 2019; Mahabadi et al., 2020; Clark et al., 2020a; Ghaddar et al., 2021; Liu et al., 2021; Creager et al., 2021; Yu et al., 2022). Moreover, the performance of those approaches on OOD data is considerably low without the hyperparameter tuning on OOD data (Yang et al., 2023).
@@ 3.1 Training Phase: Mixture Model to Capture Latent Features
We use this penalty term with the following modifications. First, the penalty cannot be minimized to zero when the mini-batch size exceeds the number of experts . During joint minimization with the classification loss , forcing the minimization of the never-zero penalty could lead models too far away from the optimal solution for . To avoid this, we consider that in the -th row of , the corresponding expert for captures the same latent features among the top- elements (instances). We then exclude those elements to allow such multi-instance assignments. Let be a function to drop out the elements with the top- values in each row to 0 (Algorithm 1). We minimize .
@@ 4 Experiments
Our goal is to achieve predictions robust to distribution shifts related to shortcuts. In this section, we test whether the proposed post-hoc control improves performance on those OOD tests and analyze the mechanism based on our assumption.
@@ 4.3 Analyses
We analyzed the contribution to the performance with respect to the hyperparameters of our mixture model: the number of experts , the number of row-wise dropouts , and the loss-weighting value . Table 4 shows the ablation study on MNLI. This table shows how performance changes by varying one of the hyperparameters from the values determined to be optimal on the ID dev. There is little to no difference in performance on the ID dev for any given value, but the OOD performance with post-hoc control is best for nearly all the values determined to be best on the ID dev. It is also worth noting that using our and top- dropout consistently improves OOD performance better than without using them (when or is zero). These results indicate the effectiveness of the proposed training strategy and hyperparameter search.
@@ 5 Related Work
Several approaches used the counterfactual framework of causal inference. To make counterfactual predictions unaffected by shortcuts, Tian et al. (2022); Niu et al. (2021) combined predictions of a main model and a shortcut-dependent model. Wang and Culotta (2020) classified features into genuine or spurious and selected genuine features for predictions. Others utilized identified spurious features to train model predictions to be invariant to interventions on the spurious features (Veitch et al., 2021; Makar et al., 2022; Puli et al., 2022b).
@@
The above methods still require OOD test data related to pre-identified shortcuts to tune hyperparameters, as described in Section 2.3. Our method is different from them in that the training and tuning can be conducted solely on ID data. In Section 4.2, we demonstrated that in the setting of fully unknown shortcuts where only ID data are available, our method improves the performance on OOD data significantly better than the previous methods.
@@
Outside of NLU, shortcuts have been addressed in the ML literature as one of the broader OOD problems (Krueger et al., 2021; Yang et al., 2023). Still, many methods used in ML and NLU tasks have the same concepts in common, such as reweighting (Nam et al., 2020; Liu et al., 2021; Clark et al., 2020a; Utama et al., 2020b), IRM (Creager et al., 2021; Yu et al., 2022), counterfactual invariance (Veitch et al., 2021; Makar et al., 2022; Puli et al., 2022b), and data augmentation (Yao et al., 2022; Puli et al., 2022a; Wu et al., 2022). As described in Section 2.2, IRM (Arjovsky et al., 2019) and GroupDRO (Sagawa et al., 2020) are the two principal approaches. These approaches considered the known shortcut setting, and similar to NLU literature, their follow-up approaches have sought to address shortcuts in the unknown shortcut setting (Nam et al., 2020; Liu et al., 2021; Creager et al., 2021; Yao et al., 2022; Puli et al., 2022a; Izmailov et al., 2022; Kirichenko et al., 2023). However, also similar to NLU literature, those follow-up approaches still require the shortcuts that shift in test data to be pre-identified in validation data (Yang et al., 2023).
@@ Appendix A Additional Figures
Appendix B Further Setup Details
Table 5 specifies the URLs of the datasets, pre-trained models, and code of the previous methods we introduced in Section 4.1.
@@ Appendix B Further Setup Details
Datasets MNLI https://cims.nyu.edu/~sbowman/multinli/ HANS https://github.com/tommccoy1/hans QQP and PAWS https://github.com/google-research-datasets/paws FEVER and Symm. v1/v2 https://github.com/TalSchuster/FeverSymmetric Pre-Trained Models BERT https://huggingface.co/bert-base-uncased DeBERTa https://huggingface.co/microsoft/deberta-v3-large Code Conf-reg ∗ https://github.com/UKPLab/emnlp2020-debiasing-unknown JTT∗ https://github.com/YyzHarry/SubpopBench RISK https://github.com/CuteyThyme/RISK EIIL https://github.com/PluviophileYU/BAI BAI https://github.com/PluviophileYU/BAI GroupDRO∗ https://github.com/YyzHarry/SubpopBench ReWeightCRT∗ https://github.com/YyzHarry/SubpopBench